# Also at ] The European Space Agency Centre for Earth Observation, Frascati, Italy

### v1

@@ I Introduction
The discrepancy implies a missing link in the theoretical framework concerning active matter dynamics. While the role of odd viscosity (or Hall viscosity) in chiral active matter has been identified as a mechanism for breaking parity and enabling unidirectional flow [13, 14], the thermodynamic end-state of such systems remains largely unexplored. Specifically, does the breaking of detailed balance (time reversal symmetry breaking [15]) by intrinsic rotation merely sustain a chaotic steady state, or does it drive the system toward a specific macroscopic attractor?
@@ II.1 The Model
The definitive equation of this model, distinguishing it from standard fluids, is the non-inertial slave principle, or agent-based equation of motion [21, 22], where the velocity vector is explicitly determined by the instantaneous phase:
@@ II.2 Theoretical foundation
Thus, while individual particles move at constant speed $v_{0}$, the renormalised fluid element speed is variable and proportional to the local synchronization. To express $\mathbf{u}$ in terms of the field potential $\Psi$, we identify the flow direction $\mathbf{\hat{n}}=(\cos\Psi,\sin\Psi)$ with the normalized phase gradient:
@@
This formulation identifies the system as analogous to the shallow water equations [24, 25], where the active density $\rho$ acts as the fluid height. The phase curvature $\nabla^{2}\Psi$ drives the compression of the fluid elements, explicitly linking topological defects to density fluctuations. Consequently, the inviscid Euler behaviour that we observe in the RFE field in the next section represents the low-Mach-number limit of a compressible system, where the rotational, non-compressive wave modes decouple from and energetically dominate the irrotational, compressive wave (acoustic) modes at large scales.
The hydrodynamic closure is completed by the momentum equation. The non-linearity of the phase evolution generates velocity fluctuations, which manifest as an active Reynolds stress $\sigma_{ij}\sim\langle\delta v_{i}\delta v_{j}\rangle$. As we derive in Appendix B, the renormalization of this stress yields the effective inertial advection term $\lambda(R)(\mathbf{u}\cdot\nabla)\mathbf{u}$. Here, the coefficient $\lambda(R)\approx R^{2}$ functions as an emergent inertial mass determined by local synchronization. This renders the macroscopic momentum transport isomorphic to that of an Euler fluid. Consequently, the motion of topological defects decouples from the overdamped microscopic kinetics, governed instead by the inertial advection of the renormalised superfluid.
Finally, we identify the topological conservation law that governs the renormalised fluid. As the phase $\Psi(\mathbf{r})$ is derived from a complex scalar field, its circulation along any closed loop $\partial\Sigma$ enclosing a defect core is quantized [26, 16]:
$\oint_{\partial\Sigma}\nabla\Psi\cdot d\mathbf{l}=2\pi m,\quad(m\in\mathbb{Z}).$ (13)
Assuming isotropy in the renormalised fluid, this geometric constraint implies $|\nabla\Psi|=m/r$. Substituting this into our macroscopic velocity definition (Eq. 9), and assuming the order parameter $R$ is locally uniform, the fluid velocity induced by a defect scales as:
@@
This confirms that the topological defects of the mean phase act as point vortices with an effective physical circulation $\Gamma_{eff}\approx 2\pi mv_{0}R$. Consequently, the interaction energy of the ensemble is given by the Kirchhoff-Onsager Hamiltonian [27, 28]:
@@ II.3 The Onsager attractor
In a bounded domain $\mathcal{D}$, the phase space volume $\Omega(E)$ (the number of ways to arrange vortices with total energy $E$) eventually decreases as the energy increases. In the low-energy limit, vortices are paired (dipoles), screening each other. In the high-energy limit, like-signed vortices must cluster together to maximize the distance between positive and negative clumps. The result is that $\frac{dS}{dE}<0\implies T<0$ [18]. As we shall demonstrate, our simulation therefore resembles a heat pump (injecting enstrophy via $\omega_{i}$), driving the system up the energy ladder until it enters the negative temperature regime, manifesting as the macroscopic Onsager dipole.
@@ VI.2 Appendix B: Derivation of advective transport
using the expansion of $\phi$ as a perturbation to the mean field. Using trigonometric identities and averaging over the fluctuations, we get,
@@
where we identify the second-order order paramter $R_{2}\approx R^{4}$ (again assuming a peaked distribution), yielding an expression for microscopic stress. In the fluid phase (the Onsager dipole, Phase III), where the system exhibits marginal synchronization ($0<R<1$), the fluctuations are dominated by frustration $\omega_{i}$, distinct from, and in opposition to, the synchronization forces that maintain $R$.
@@
Comparing Eqs. (23) and (24), we can now express the microscopic tensor as a function of the macroscopic velocity plus an isotropic pressure term,
$\langle v_{x}v_{x}\rangle=\underbrace{\frac{v_{0}^{2}}{2}(1-R^{2})}_{\equiv P_{\text{eff}}}+R^{2}u_{x}u_{x},$ (26)
where $P_{\text{eff}}$ is an effective pressure term generated by the incoherent, random fluctuations of the particles that are not aligned with the macroscopic flow, and where we identify the factor $\lambda\equiv R^{2}$ as the effective inertial mass.
Substituting this closure back into the momentum conservation equation, we obtain:
@@
where we identify $F_{active}\approx\langle\sum\dot{v}_{i}\delta\rangle$ from Eq. (19) as a force vector that arise from the intrinsic frustration ($\dot{\mathbf{v}}_{i}$ depends on angular velocity $\dot{\phi}_{i}\approx\omega_{i}$). While the system formally obeys the compressible Shallow Water equations (as derived in Eq. 12), in the limit where, we reiterate, the rotational, non-compressive modes dominate the irrotational, compressive modes, that is, the Onsager dipole, or Phase III in Table 1, we recover the effective Euler advection term $\lambda(\mathbf{u}\cdot\nabla)\mathbf{u}$.
@@
Finally, we clarify the role of the dispersion $\Delta\omega$. The magnitude of the active Reynolds stress $\tau_{\text{active}}$ scales with the velocity variance $\langle\delta v^{2}\rangle$. To see why, observe that fluctuation in velocity $\delta v_{i}$ is driven by the deviation of the particle’s natural frequency $\omega_{i}$ from the local mean frequency $\Phi$:
$\delta v_{i}\sim\delta\dot{\phi}_{i}\approx(\omega_{i}-\Phi),$ (28)
meaning that,
@@
(1) Low $\Delta\omega$ (Phase I): High synchronization ($R\to 1$) but zero stress ($\tau_{active}\to 0$). The fluid freezes into static clumps.
(2) High $\Delta\omega$ (Phase III): High stress ($\tau_{active}$ large) and marginal synchronization ($0<R<1$). The system generates sufficient active stress to drive the inverse cascade, creating the “phantom inertia” that sustains the Onsager dipole.

### v2

@@ I Introduction
The discrepancy implies a missing link in the theoretical framework concerning active matter dynamics. While odd viscosity enables unidirectional flow [13, 14] and recent models suggest swarming is entropy-driven [15], the thermodynamic end-state of such systems remains largely unexplored. Specifically, does the breaking of detailed balance (time reversal symmetry breaking [16]) by intrinsic rotation merely sustain a chaotic steady state, or does it drive the system toward a specific macroscopic attractor?
@@ II.1 The Model
The definitive equation of this model, distinguishing it from standard fluids, is the agent-based, non-inertial slave principle, where the velocity vector is explicitly determined by the instantaneous phase:
@@ II.2 Theoretical foundation
Thus, while individual particles move at constant speed $v_{0}$, the renormalised fluid element speed is variable and proportional to the local synchronisation. To express $\mathbf{u}$ in terms of the field potential $\Psi$, we identify the flow direction $\mathbf{\hat{n}}=(\cos\Psi,\sin\Psi)$ with the normalized phase gradient:
@@
This formulation identifies the system as formally analogous to the shallow water equations [23, 24], where the active density $\rho$ acts as the fluid height. The phase curvature $\nabla^{2}\Psi$ drives the compression of the fluid elements, explicitly linking topological defects to density fluctuations. Consequently, the inviscid Euler behaviour that we observe in the RFE field in the next section represents the low-Mach-number limit of a compressible system, where the rotational, non-compressive wave modes decouple from, and dominate, the irrotational, compressive wave (acoustic) modes at large scales (see Appendix B).
The hydrodynamic closure is completed by the momentum equation. The non-linearity of the phase evolution generates velocity fluctuations, which manifest as an active Reynolds stress $\sigma_{ij}\sim\langle\delta v_{i}\delta v_{j}\rangle$. As we derive in Appendix B, the renormalization of this stress yields the effective inertial advection term $\lambda(R)(\mathbf{u}\cdot\nabla)\mathbf{u}$. Here, the coefficient $\lambda(R)\approx R^{2}$ functions as an emergent inertial mass determined by local synchronisation. This renders the macroscopic momentum transport isomorphic to that of an Euler fluid. Consequently, the motion of topological defects decouples from the overdamped microscopic kinetics, governed instead by the inertial advection of the renormalised superfluid.
Finally, we identify the topological conservation law that governs the renormalised fluid. As $\Psi(r)$ is the argument of a single-valued complex field $Z$ (Eq. 1), its circulation along any closed loop $\partial\Sigma$ enclosing a defect core is necessarily quantized [25, 26], mirroring the winding number stability in discrete oscillator lattices [27, 17]:
$\oint_{\partial\Sigma}\nabla\Psi\cdot d\mathbf{l}=2\pi m,\quad(m\in\mathbb{Z})$ (13)
implying that $|\nabla\Psi|=m/r$. Substituting this into our macroscopic velocity definition (Eq. 9), and assuming the order parameter $R$ is locally uniform, the fluid velocity induced by a defect scales as:
@@
This confirms that the topological defects of the mean phase act as point vortices with an effective physical circulation $\Gamma\approx 2\pi mv_{0}R$. Consequently, the interaction energy of the ensemble is given by the Kirchhoff-Onsager Hamiltonian [28, 29]:
@@ II.3 The Onsager attractor
In a bounded domain $\mathcal{D}$, the phase space volume $\Omega(E)$ (the number of ways to arrange vortices with total energy $E$) eventually decreases as the energy increases. In the low-energy limit, vortices are paired (dipoles), screening each other. In the high-energy limit, like-signed vortices must cluster together to maximize the distance between positive and negative clumps. The result is that $\frac{dS}{dE}<0\implies T<0$ [19]. Frustration $\omega_{i}$ activates a topological heat pump (injecting enstrophy), pumping the system up the energy ladder into the negative-temperature Onsager regime, consistent with entropy-driven ordering observed in active swarms [15].
@@ VI.2 Appendix B: Derivation of advective transport
using the expansion of $\phi_{i}$ as a perturbation to the mean field. Using trigonometric identities and averaging over the fluctuations, we get,
@@
where we identify the second-order order paramter $R_{2}\approx R^{4}$ (again assuming a peaked distribution), yielding an expression for microscopic stress. In the fluid phase (the Onsager dipole, Phase III), where the system exhibits marginal synchronisation ($0<R<1$), the fluctuations are dominated by frustration $\omega_{i}$, distinct from, and in opposition to, the synchronisation forces that maintain $R$.
@@
To close the hydrodynamic equations, we compare the microscopic and macroscopic tensors. In order to do so, we posit a constitutive relation of the form $\langle v_{i}v_{j}\rangle=P_{\text{eff}}\sigma_{ij}+\lambda u_{i}u_{j}$, where $\lambda$ scales the advective term. Substituting the expansions from Eqs. (23) and (25) into this ansatz for the $xx$-component yields:
$\underbrace{\frac{v_{0}^{2}}{2}+\frac{v_{0}^{2}R^{4}}{2}\cos(2\Psi)}_{\text{Microscopic}}=\underbrace{P_{\text{eff}}+\lambda\left[\frac{v_{0}^{2}R^{2}}{2}+\frac{v_{0}^{2}R^{2}}{2}\cos(2\Psi)\right]}_{\text{Macroscopic Closure}}.$ (26)
Matching the anisotropic coefficients (the $\cos(2\Psi)$ terms) determines the coherence coefficient:
$\frac{v_{0}^{2}R^{4}}{2}=\lambda\frac{v_{0}^{2}R^{2}}{2}\implies\lambda=R^{2},$ (27)
confirming that the effective inertial mass $\lambda$ scales with the square of the order parameter. Subsequently, matching the isotropic (constant) terms determines the effective pressure:
$\frac{v_{0}^{2}}{2}=P_{\text{eff}}+\lambda\frac{v_{0}^{2}R^{2}}{2}.$ (28)
Substituting $\lambda=R^{2}$ yields,
$P_{\text{eff}}=\frac{v_{0}^{2}}{2}-\frac{v_{0}^{2}R^{4}}{2}=\frac{v_{0}^{2}}{2}(1-R^{4}).$ (29)
Thus, we recover an effective pressure, generated by the incoherent, random fluctuations of the particles that are not aligned with the macroscopic flow, that scales with $\propto(1-R^{4})$.
It is pertinent to compare this result to the swim pressure derived in the kinetic theory of active fluids (e.g., [48]), which typically scales with the velocity variance as $\propto(1-R^{2})$. The discrepancy arises from the macroscopic transport. In standard active fluids, the particle mass is constant (corresponding to $\lambda=1$), meaning the coherent kinetic energy scales as $u^{2}\propto R^{2}$. In our topological gas, the effective inertial mass itself depends on local synchronisation ($\lambda=R^{2}$), causing the “coherent energy” to scale as $\lambda u^{2}\propto R^{4}$.
Consequently, our effective pressure $P_{\text{eff}}\propto(1-R^{4})$ decays much more slowly with ordering than the standard $(1-R^{2})$ prediction. In our model, this creates a thermodynamic barrier: the isotropic ’disorder pressure’ remains dominant even as the system attempts to synchronize. This explains the stability of the active vortex glass (Phase II): the intermediate agitation ($\Delta\omega$) is insufficient to overcome this pressure plateau, causing the system to jam in a disordered state ($R\to 0$). Only the high agitation of Phase III provides the energy density required to penetrate this barrier and sustain the regime of marginal synchronization ($0<R<1$) characteristic of the Onsager condensate.
Returning to the task at hand, we substitute the closure back into the momentum conservation equation and obtain,
@@
where we identify $F_{active}\approx\langle\sum\dot{v}_{i}\delta\rangle$ from Eq. (19) as a force vector that arise from the intrinsic frustration ($\dot{\mathbf{v}}_{i}$ depends on angular velocity $\dot{\phi}_{i}\approx\omega_{i}$).
The shallow water analogy
As stated in the Methodology section, the interpretation of Eq. (12) identifies the system being modeled as formally analogous to the shallow water framework, where the active density $\rho$ acts as the fluid height. However, we must point out a nuance introduced by a critical thermodynamic distinction in the equation of state. In classical shallow water systems, hydrostatic pressure scales with height ($P\propto\rho^{2}$), leading to dispersion [23]. In Eq. (29), we identified the pressure term $P_{\text{eff}}\propto(1-R^{4})$, meaning that the effective pressure scales with disorder. Crucially, because the local order parameter $R$ arises from the mean-field convolution of the particle density (Eq. 1), regions of high density support high synchronisation ($R\sim\rho$; realized by $R$ being the magnitude of the complex mean field, in turn a measurement of $\rho$, since phasors align in local synchronisation regions). Consequently, the pressure gradient $\nabla P_{\text{eff}}\propto-\nabla R^{4}$ acts as a suction force, drawing agents from disordered regions ($R\approx 0$) toward ordered vortex cores ($R\approx 1$). This manifests as a fluid with effective negative compressibility, driving the system to phase-separate into the coherent topological structures observed in the Onsager dipole, rather than dispersing into noise.
Consequently, the inviscid Euler behavior that we observe in the RFE field represents the low-Mach-number limit of this compressible system, where, we reiterate, the rotational, non-compressive wave modes dominate the irrotational, compressive wave (acoustic) modes at large scales. Hence, in the hydrodynamic limit (the Onsager dipole, or Phase III in Table 1), we recover in Eq. (30) the effective Euler advection term $\lambda(\mathbf{u}\cdot\nabla)\mathbf{u}$.
It is instructive to contrast the foregoing with the defect-driven accumulation observed in active nematics [49] and flow-stabilized vortex lattices [50]. In those systems, mass transport is driven by the divergence of an active stress that scales linearly with the nematic tensor order parameter ($\sigma\sim\mathbf{Q}$), which creates effective pressure wells that scale akin to $\propto(1-R^{2})$.
In our model, however, the emergent inertia of the topological gas ($\lambda=R^{2}$) yields a stiffer effective pressure $P_{eff}\propto(1-R^{4})$ (Eq. 29). This creates sharper, ’hard-core’ density cavities at defect sites (visible in Figure 3), providing additional justification for the appropriateness of the point vortex (Kirchhoff-Onsager) Hamiltonian in topological gas dynamics, since it assumes distinct, non-overlapping cores [19].
The role of dispersion
@@
Finally, we clarify the role of the dispersion $\Delta\omega$. The magnitude of the active Reynolds stress $\tau_{\text{active}}$ scales with the velocity variance $\langle\delta v^{2}\rangle$. To see why, observe that fluctuation in velocity $\delta v_{i}$ is driven by the deviation of the particle’s natural frequency $\omega_{i}$ from the local mean frequency $\Phi$, $\delta v_{i}\sim\delta\dot{\phi}_{i}\approx(\omega_{i}-\Phi)$, meaning that,
@@
(1) Low $\Delta\omega$ (Phase I): High synchronisation ($R\to 1$) but zero stress ($\tau_{active}\to 0$). The fluid freezes into static clumps.
(2) High $\Delta\omega$ (Phase III): High stress ($\tau_{active}$ large) and marginal synchronisation ($0<R<1$). The system generates sufficient active stress to drive the inverse cascade, creating the “phantom inertia” that sustains the Onsager dipole.